### v5

# LayerAct: Advanced Activation Mechanism for Robust Inference of CNNs

Abstract
In this work, we propose a novel activation mechanism called LayerAct for CNNs. This approach is motivated by our theoretical and experimental analyses, which demonstrate that Layer Normalization (LN) can mitigate a limitation of existing activation functions regarding noise robustness. However, LN is known to be disadvantageous in CNNs due to its tendency to make activation outputs homogeneous. The proposed method is designed to be more robust than existing activation functions by reducing the upper bound of influence caused by input shifts without inheriting LN’s limitation. We provide analyses and experiments showing that LayerAct functions exhibit superior robustness compared to ElementAct functions. Experimental results on three clean and noisy benchmark datasets for image classification tasks indicate that LayerAct functions outperform other activation functions in handling noisy datasets while achieving superior performance on clean datasets in most cases.
Code & Appendixhttps://github.com/KihyukYoon/LayerAct
Introduction
Most existing activation functions follow element-level activation (ElementAct) mechanisms, meaning the functions apply independently to each element of the input. However, we identified a limitation in the method: their robustness varies across samples. This is because their robustness relies on the saturation state, where the activation output for a specific range converges to a certain value. For example, in a layer with ReLU (Nair and Hinton 2010), the elements of the output remain unaffected by input shifts only when they are in the saturation state, specifically when the input and shifted input of the elements are smaller than zero. Consequently, existing activation functions can ensure robustness only when a sufficient number of elements are in the saturation state, not when there are fewer elements in that state.
We found that Layer Normalization (LN; Ba, Kiros, and Hinton (2016)) can address this limitation of ElementAct functions. Our theoretical and experimental analyses reveal that placing LN right before the activation (i.e., using the output of LN as the input to the activation) can enhance the robustness of networks. However, previous studies have reported that Convolutional Neural Networks (CNNs) with LN tend to perform poorly on clean datasets, mainly due to LN’s tendency to lead to homogeneous activation. These claims are supported by our empirical results on the representation of the last layer in CNNs with LN. Therefore, our goal is to develop a method that captures the noise-robustness benefits of LN without inheriting its homogeneity limitations.
To achieve this goal, we propose a novel activation method, the Layer-level Activation (LayerAct) mechanism. Unlike the ElementAct mechanism, our proposed method utilizes layer-direction normalization for activation, but the activation output is far different from that of ElementAct with LN. This enables LayerAct to absorb LN’s robustness benefit for activation while avoiding its homogeneity limitation. Additionally, LayerAct functions are not subject to the trade-off issue between two important properties of activation, one-side saturation and zero-like activation mean (Clevert, Unterthiner, and Hochreiter 2016; Qiu, Xu, and Cai 2018), which are faced by ElementAct functions (for details, see Appendix A). These two benefits of LayerAct functions, enhanced robustness and freedom from the trade-off problem, underscore their potential to outperform existing ElementAct functions on both noisy and clean datasets.
Experimental analysis with the MNIST dataset demonstrates that LayerAct functions have the following properties: (1) the mean activation of LayerAct functions is zero-like, and (2) the output fluctuation due to noisy input is smaller with these functions than with ElementAct functions. We compared the performance of the residual networks (ResNets; He et al. (2016)) with LayerAct functions to those with other ElementAct functions on three image classification tasks.
Our contributions can be summarized as follows:
  • We identify a previously unrecognized limitation in existing activation mechanisms, namely a large variance of noise-robustness across samples.
  • We present analysis and experimental evidence showing that while LN can address this limitation, it also causes the final layer representation in CNNs to be less distinguishable between samples with different labels.
  • We introduce LayerAct that exhibit greater robustness than existing ElementAct functions, without the homogeneity limitation of LN.
  • We empirically evaluate LayerAct functions, and they show superior performance on both clean and noisy datasets in most cases.
Backgrounds
In this section, we provide the background for the analyses presented in the remainder of this paper, including the definition of activation scale and a review of normalizations.
Activation Scale
In some activation functions, a function bounded between one and zero characterizes their non-linearity. We define this function, denoted as $s$, and its output as the activation scale function and activation scale, respectively.
Consider a layer in a multi-layer perceptron with linear projection and an activation function. Given a $r$-dimensional vector $x=(x_{1},...,x_{r})^{T}$, a weight matrix $W\in\mathbb{R}^{r\times D}$, where d is the dimension of activation input (e.g., $D=C\times H\times W$ for image), and a non-linear activation function $f$, the activation output is $a=f\left(y\right)$, where $y=W^{T}x$. With an activation scale function $s$, the activation output and gradient of forward and backward passes are:
$a_{i}=y_{i}s\left(y_{i}\right),\quad\frac{\partial a_{i}}{\partial y_{i}}=s\left(y_{i}\right)+y_{i}\frac{\partial s\left(y_{i}\right)}{\partial y_{i}},$ (1)
where $s$ is increasing and $s\left(y_{i}\right)>0$ if $y_{i}>0$. For example, the activation scale function for the $i^{th}$ element in the SiLU (Elfwing, Uchibe, and Doya 2018) is presented as follows:
$\begin{split}&s^{SiLU}\left(y_{i}\right)=sigmoid\left(y_{i}\right)=\frac{1}{1+e^{-y_{i}}},\end{split}$ (2)
where $sigmoid$ and $s^{SiLU}$ represent the Logistic Sigmoid function, and the activation scale functions of SiLU, respectively. For further discussion, we define a specific activation scale function $s$.
Definition 1
(Activation Scale Function) The activation scale function $s$ is an increasing Lipschitz continuous function that is bounded between zero and one:
$s\left(0\right)=1/2,\;\left|s\left(a\right)-s\left(b\right)\right|\leq K\left|a-b\right|\quad\forall a,b\in\mathbb{R}.$ (3)
For example, the activation scale of SiLU satisfies this, as the activation scale is $sigmoid\left(y_{i}\right)$. Meanwhile, the activation scale also determines the saturation state of the activation functions, which occurs when $s\left(y_{i}\right)\simeq 0$.
In conclusion, the activation scale function plays a crucial role in providing non-linearity during the forward pass, controlling the gradient during the backward pass, and determining the saturation state of the activation function.
Revisiting Normalizations in CNNs
Normalization layers re-scale and re-center both the input and the gradient during forward and backward propagation, respectively. The normalization output $n_{i}$ from the normalization input $y_{i}$ is:
$\begin{split}n_{i}=\gamma_{i}\cdot\frac{y_{i}-\mu_{y}}{\sigma_{y}}+\beta_{i},\end{split}$ (4)
where $\gamma_{i}$ and $\beta_{i}$ are the learnable parameters of the affine function, $D$ is the normalizing dimension, and $\mu_{y}$ and $\sigma_{y}$ are:
$\begin{split}\mu_{y}=\frac{1}{D}\sum^{D}_{i=1}y_{i},\quad\sigma_{y}=\sqrt{\frac{1}{D}\sum^{D}_{i=1}\left(y_{i}-\mu_{y}\right)^{2}},\end{split}$ (5)
the mean and variance of the input $y$, respectively. Normalization layers can stabilize the network training by re-scaling and re-centering as shown in Equation 4. The dimension $D$ varies according to the normalization methods, $D=N\times H\times W$ for Batch Normalization (BN; Ioffe and Szegedy (2015)) and $D=C\times H\times W$ for LN, where $N$, $C$, $H$, and $W$ denote the size of the batch size, channel, height, and width of the image dataset.
BN has achieved great success in computer vision tasks within CNN-based networks and remains dominant in the field. In contrast, LN is less preferred compared to BN, largely due to the inferior performance of networks using LN. Previous studies have attributed this underperformance to LN’s tendency to produce homogeneous outputs (Lubana, Dick, and Tanaka 2021; Labatie et al. 2021).
Problems Analysis and Research Motivation
In this section, we present our motivation, specifically, we focus on: (1) an analysis of the limitations of existing activation functions, and (2) theoretical and empirical analyses of the advantages and disadvantages of LN in CNNs.
Large Variance of Activation Robustness
To analyze the robustness, we define activation fluctuation that can represent the influence of the shift of inputs.
Definition 2
(Activation Fluctuation) Let $\epsilon=(\epsilon_{1},\epsilon_{2},...,$ $\epsilon_{D})^{T}$ be the input shift vector, which is independent of the input vector $y$. We define activation fluctuation as $\left\|f\left(y+\epsilon\right)-f\left(y\right)\right\|\leq c$, where the constant $c$ is the upper bound of activation fluctuation.
The lower the upper bound $c$ is, the lower the variance of robustness across samples. We can define the activation fluctuation of ElementAct functions.
Definition 3
(Activation Fluctuation of Element-level  Activation Functions) Let $\hat{y}_{i}=y_{i}+\epsilon_{i}$ be the input with a shift. The activation fluctuation of an ElementAct function $f$ is given by:
$\begin{split}\left\|f\left(\hat{y}\right)-f\left(y\right)\right\|&=\sum^{D}_{i=1}\left|\hat{y}_{i}s\left(\hat{y}_{i}\right)-y_{i}s\left(y_{i}\right)\right|\\ &=\sum^{D}_{i=1}\left|y_{i}\left(s\left(\hat{y}_{i}\right)-s\left(y_{i}\right)\right)+\epsilon_{i}s\left(\hat{y}_{i}\right)\right|.\end{split}$ (6)
A sample will exhibit a small $\left\|f\left(\hat{y}\right)-f\left(y\right)\right\|$ if a sufficient number of its elements are in a saturation state. However, ElementAct functions do not ensure that all samples have a sufficient number of elements in the saturation state. More specifically, the activation fluctuation is upper-bounded when not all elements are in the saturation state, where $y_{i}>0$ for all $i$:
$\begin{split}\left\|f\left(\hat{y}\right)-f\left(y\right)\right\|\leq\sum^{D}_{i=1}\left(y_{i}\left|s\left(\hat{y}_{i}\right)-s\left(y_{i}\right)\right|+\left|\epsilon_{i}\right|\cdot s\left(\hat{y}_{i}\right)\right).\end{split}$ (7)
Considering Definition 1, the upper bound of $\left\|s\left(\hat{y}\right)-s\left(y\right)\right\|$ and $\left\|s\left(\hat{y}\right)\right\|$ of ElementAct functions are given by the following, respectively:
$\left\|s\left(\hat{y}\right)-s\left(y\right)\right\|\leq\sum^{D}_{i=1}K\left|\epsilon_{i}\right|,\quad\left\|s\left(\hat{y}_{i}\right)\right\|\leq d.$ (8)
Equation 7 demonstrates that activation scale is closely related to the activation fluctuation. Samples with large $\left\|s\left(\hat{y}\right)-s\left(y\right)\right\|$ and $\left\|s\left(\hat{y}\right)\right\|$ are not robust to the input shift $\epsilon$. Thus, a method that has a lower upper bound for $\left\|s\left(\hat{*}\right)-s\left(*\right)\right\|$ and $\left\|s\left(\hat{*}\right)\right\|$ will reduce the upper bound of activation fluctuation, resulting in a low variance of robustness across samples.
Empirical Analysis on Normalizations in CNNs
To empirically analyze the impact of BN and LN in CNNs, we conducted a series of experiments using the CIFAR10 (Krizhevsky 2009) dataset with the ResNet20 architecture. In the original ResNet20, BN layers are positioned immediately before the activation layers. For our experiments, we designed four variants of ResNet20: (1) eliminating BN layers entirely (ResNet20-None), (2) replacing BN layers with LN layers (ResNet20-LN), (3) adding LN layers before the BN layers (ResNet20-(LN, BN)), and (4) adding LN layers after the BN layers (ResNet20-(BN, LN)).
Data BN None LN BN, LN LN, BN
C10 91.29 88.51 88.24 90.65 89.73
C10-C 69.92 67.43 74.5 74.34 72.55
C10 91.45 88.29 87.53 89.74 90.17
C10-C 70.12 68.94 74.96 73.69 72.47
Table 1: Classification performance of original ResNet20 and three variants of ResNet20 with ReLU (upper) and SiLU (lower) on CIFAR10 (C10) and CIFAR10-C (C10-C).
Table 1 demonstrates the performance of the ResNet20 variants with different normalization on both CIFAR10 and CIFAR10-C (Hendrycks and Dietterich 2019). The networks were trained on CIFAR10. The CIFAR10-C dataset serves as a noise-robustness benchmark, incorporating 19 out-of-distribution (OOD) corruptions of CIFAR10. This dataset includes a total of 19 distinct corruptions, each with five levels of severity, organized into five categories: noise, blur, digital, weather, and extra. We evaluated the models using two activation functions: ReLU and SiLU. The table reports the average mean accuracy over 30 runs, with ”Norm” and ”Act” denoting the normalization and activation, respectively. The experimental results show that the use of LN alone leads to inferior performance on the clean dataset, CIFAR10. Specifically, networks with LN alone showed inferior results compared to those without normalization, which aligns with findings from previous studies.
LN and other batch-free normalizations have been shown to block the negative effect of BN on robustness of networks (huang2022delving). Interestingly, our experiments reveal a detailed robustness advantage of LN, which can lead to robust activation, and this has not been emphasized in previous studies. On the noisy dataset, CIFAR10-C, the networks incorporating LN demonstrated better performance than those without LN. Notably, when LN is applied immediately before the activation layers, as in ResNet20-LN and ResNet20-(BN, LN), these networks outperformed their counterparts, ResNet20-None and ResNet20-(LN, BN), on noisy datasets, despite underperforming on the clean dataset. This suggests that LN may provide a robustness benefit, which is closely related to activation.
LN and Activation Robustness
Based on the experimental result from the experimental analysis, we analyzed the relationship between LN and activation robustness. Here, we only consider LN without an affine function. In this case, the normalized output $n_{i}$, the normalized output with a shift $\hat{n}_{i}$, the activation output $a$, and activation output with shift $\hat{a}$ are as follows:
$\begin{split}n_{i}=\frac{y_{i}-\mu_{y}}{\sqrt{\sigma^{2}_{y}+\alpha}},\quad\hat{n}_{i}=\frac{\hat{y}_{i}-\hat{\mu}_{y}}{\sqrt{\hat{\sigma}^{2}_{y}+\alpha}},\\ a_{i}=n_{i}s\left(n_{i}\right),\quad\hat{a}_{i}=\hat{n}_{i}s\left(\hat{n}_{i}\right),\end{split}$ (9)
where $\hat{\mu}_{y}$ and $\hat{\sigma}_{y}^{2}$ are the layer-direction mean and variance of $\hat{y}$. The activation fluctuation using Equations 4 and 9 are:
Definition 4
(Activation Fluctuation of ElementAct Functions with LN) Let $\epsilon_{i}$, $\hat{y}_{i}$, and $\hat{n}_{i}$ of the $i^{th}$ normalization layer be the shift of the input, the input with the shift, and the output from the input. The activation fluctuation of an ElementAct function $f$ is:
$\begin{split}\left\|f\left(\hat{n}\right)-f\left(n\right)\right\|=&\sum^{D}_{i=1}\left|\hat{n}_{i}s\left(\hat{n}_{i}\right)-n_{i}s\left(n_{i}\right)\right|\allowdisplaybreaks\\ \lesssim&\sum^{D}_{i=1}\left(\left|n_{i}\right|\cdot\left|s\left(\hat{n}_{i}\right)-s\left(n_{i}\right)\right|\right)\allowdisplaybreaks\\ &+\sum^{D}_{i=1}\left(\frac{\left|\epsilon-\mu_{\epsilon}\right|\cdot s\left(\hat{n}_{i}\right)}{\sqrt{\sigma_{y}^{2}+\alpha}}\right),\end{split}$ (10)
where $\sqrt{\sigma_{y}^{2}+\alpha+\sigma_{\epsilon}^{2}}\approx\sqrt{\sigma_{y}^{2}+\alpha}$ and $\sigma_{y}\gg\sigma_{\epsilon}$.
For the detailed derivation process of the equations, see Appendix B. In the previous section, we discussed that reducing the magnitudes of the terms of the activation scale effectively decreases the upper bound of activation fluctuation, thereby reducing the variance of activation robustness across samples. Considering Definitions 1 and 4, the upper bound of $\left\|s\left(\hat{n}\right)-s\left(n\right)\right\|$ and $\left\|s\left(\hat{n}\right)\right\|$ is given as follows:
$\left\|s\left(\hat{n}\right)-s\left(n\right)\right\|<\sum^{D}_{i=1}\frac{K\left|\epsilon_{i}-\mu_{\epsilon}\right|}{\sqrt{\sigma_{y}^{2}+\alpha}},\quad\left\|s\left(\hat{n}_{i}\right)\right\|\ll d,$ (11)
which are lower than those of ElementAct functions without normalization (see Equation 8), specifically considering that $\frac{\left|\epsilon_{i}-\mu_{\epsilon}\right|}{\sqrt{\sigma_{y}^{2}+\alpha}}<1$, where $\sigma_{y}\gg\sigma_{\epsilon}$. This reveals that LN can improve activation robustness not only by re-scaling and re-centering, but also by affecting the activation scale. The experimental results in Table 1 support our analysis, demonstrating that networks with LN are much more robust compared to those without normalization.
Figure 1: Plots of the last layer representation in ResNet20 with BN (left), and LN (right).
Homogeneity Limitation of LN across Labels
It is well-known that LN produces homogeneous outputs across samples by normalizing the mean and variance of the input. However, this homogeneity extends beyond its output: LN causes activations to become similar across samples. Specifically, when LN is placed right before the activation function, producing outputs $ns\left(n\right)$, it loses the mean and variance statistics of input $y$. Here, we provide experimental results showing that LN produces homogeneous representations across samples, even those with different labels.
Figure 1 illustrates the last layer representation in ResNet20 with BN and LN on CIFAR10, visualized using Isomap embedding (Tenenbaum, de Silva, and Langford 2000). The figure demonstrates that the representation of the network with LN is less distinguishable between data points with different labels compared to the network with BN, which provides evidence of how LN leads to inferior performance on clean datasets. This reduced distinguishability is due to the similar output statistics of LN across samples. The variances of the output mean and variance in networks with LN are 0.0001 and 0.00072, respectively, which are substantially lower than those observed with BN (0.00073 and 0.00803, respectively). These experimental results align with previous studies, suggesting that LN’s tendency to produce homogeneous outputs is a key limitation.
Summary
In this section, we presented analyses and empirical results highlighting that existing activation functions have a limitation regarding noise robustness, which LN can address. However, LN is not a preferred normalization method for CNNs due to its homogeneity limitation. Motivated by this, our goal is to develop a method that captures the robustness benefit of LN without inheriting its limitation.
Figure 2: The mechanisms of the ElementAct (left) and proposed layer-level activation (right).
Layer-level Activation
In this section, we introduce LayerAct mechanism, a novel approach that operates differently from existing ElementAct functions, as illustrated in Figure 2.
LayerAct Mechanism
A function that follows LayerAct mechanism is defined as the product of the input $y_{i}$ and the activation scale $s(n_{i})$, which uses the layer-normalized input $n_{i}$. The forward and backward pass of a LayerAct function are given by:
$a_{i}=y_{i}s\left(n_{i}\right),\quad n_{i}=\frac{\left(y_{i}-\mu_{y}\right)}{\sqrt{\sigma_{y}^{2}+\alpha}},$ (12)
$\begin{split}\frac{\partial\mathcal{L}}{\partial\mu_{y}}=&\sum^{D}_{i=1}\frac{\partial\mathcal{L}}{\partial a_{i}}\cdot\frac{\partial s\left(n_{i}\right)}{\partial n_{i}}\cdot\frac{-y_{i}}{\sqrt{\sigma_{y}^{2}+\alpha}},\\ \frac{\partial\mathcal{L}}{\partial\sigma_{y}^{2}}=&\sum^{D}_{i=1}\frac{\partial\mathcal{L}}{\partial a_{i}}\cdot\frac{\partial s\left(n_{i}\right)}{\partial n_{i}}\cdot\frac{-y_{i}\cdot n_{i}}{2\left(\sigma_{y}^{2}+\alpha\right)},\\ \frac{\partial\mathcal{L}}{\partial y_{i}}=&\frac{\partial\mathcal{L}}{\partial a_{i}}s\left(n_{i}\right)+\frac{\partial\mathcal{L}}{\partial a_{i}}\cdot\frac{\partial s\left(n_{i}\right)}{\partial n_{i}}\cdot\frac{y_{i}}{\sqrt{\sigma_{y}^{2}+\alpha}}\\ &+\frac{1}{D}\cdot\frac{\partial\mathcal{L}}{\partial\mu_{y}}+\frac{2\left(y_{i}-\mu_{y}\right)}{D}\cdot\frac{\partial\mathcal{L}}{\partial\sigma_{y}^{2}},\end{split}$ (13)
respectively, where $\alpha>0$ is a constant for stability, and $\mu_{y}$ and $\sigma_{y}^{2}$ are layer-direction mean and variance, respectively.
For the stability of learning and inference, ensuring the continuity of activation outputs across the entire space is essential. While traditional ElementAct functions, such as ReLU, LReLU (Maas et al. 2013), and PReLU (He et al. 2015), do not require a continuous activation scale at zero (given that the activation output $y_{i}s\left(y_{i}\right)$ remains continuous at zero), LayerAct functions need special consideration, as the activation output $y_{i}s\left(n_{i}\right)$ is discontinuous if the activation scale function is not continuous.
Given such requirements for the activation scale functions, functions that satisfy Definition 1 are suitable for use as LayerAct functions. In this work, we propose two basic LayerAct functions, LA-SiLU and LA-HardSiLU, which utilize the Sigmoid and HardSigmoid functions as activation scale functions, respectively. The activation outputs of these functions are as follows:
$\begin{split}a^{LA\mbox{-}SiLU}=&\frac{y_{i}}{1+e^{-n_{i}}},\\ a^{LA\mbox{-}HardSiLU}=&\textit{min}\left(y_{i},\textit{max}\left(\frac{y_{i}\cdot n_{i}}{6}+\frac{y_{i}}{2},0\right)\right).\end{split}$ (14)
For the HardSigmoid of LA-HardSiLU, we used the function of Howard et al. (2019), which is a good approximation of Sigmoid.
Benefits and Properties of LayerAct
Diverse Activation Output
We discussed that LN tends to produce homogeneous activation outputs, as the mean and variance of $n_{i}$ are similar across all samples. In contrast, LayerAct functions are designed to preserve these diverse representations by transferring the statistics of the activation input $y$ to the activation output $a$. This difference is from the different activation outputs of activation with LN and LayerAct, $n_{i}s\left(n_{i}\right)$ and $y_{i}s\left(n_{i}\right)$, respectively. Due its unique output, LayerAct can maintain the input statistics within the activation outputs, resulting in much more diverse representations. For details, see Appendices C and D.
Noise-robustness of LayerAct
From here, we begin by establishing that the activation fluctuation of LayerAct is also related to the two terms of the activation scale function, $\left\|s\left(\hat{*}\right)-s\left(*\right)\right\|$ and $\left\|s\left(\hat{*}\right)\right\|$, as outlined in previous sections. Subsequently, we show that these two terms for LayerAct are bound to be lower than those of ElementAct. We can define the activation fluctuation of LayerAct as follows.
Definition 5
(Activation Fluctuation of LayerAct Functions) The activation fluctuation of a LayerAct activation function $g$, where $\hat{n}_{i}=\left(\hat{y}_{i}-\mu_{\hat{y}}\right)/\sigma_{\hat{y}}$ denotes the $i^{th}$ noisy normalized input, is defined as:
$\begin{split}\left\|g\left(\hat{y}\right)-g\left(y\right)\right\|=&\sum^{D}_{i=1}\left|\hat{y}_{i}s\left(\hat{n}_{i}\right)-y_{i}s\left(n_{i}\right)\right|\\ =&\sum^{D}_{i=1}\left|y_{i}\left(s\left(\hat{n}_{i}\right)-s\left(n_{i}\right)\right)+\epsilon_{i}s\left(\hat{n}_{i}\right)\right|.\end{split}$ (15)
This definition enables us to establish an upper bound for the activation fluctuation of LayerAct functions as follows:
$\left\|g\left(\hat{y}\right)-g\left(y\right)\right\|\leq\sum^{D}_{i=1}\left(\left|y_{i}\right|\left|s\left(\hat{n}_{i}\right)-s\left(n_{i}\right)\right|+\left|\epsilon_{i}\right|s\left(\hat{n}_{i}\right)\right).$ (16)
The terms of LayerAct’s activation scale functions, $\left\|s\left(\hat{n}\right)-s\left(n\right)\right\|$ and $\left\|s\left(\hat{n}\right)\right\|$, align precisely with those of the activation with LN, as indicated in Equation 11. Comparing Equations 8 and 11 reveals that LayerAct functions have a lower upper bound of activation fluctuation, similar to activation with LN. This reduction is achieved without the direct re-centering and re-scaling of the activation output. Therefore, it suggests that networks with LayerAct are likely to exhibit improved robustness during the forward pass.
Addressing the Trade-off
LayerAct can address another limitation of existing activation functions, namely a trade-off between two important properties of activation: one-side saturation and zero-like activation mean. The key distinction in saturation between the element-level and LayerAct functions is that element-level functions have a fixed saturation state at a certain point of activation output, whereas the saturation state of LayerAct functions depends on layer-dimension normalized inputs. Consequently, LayerAct functions can still reach a saturation state where $s\left(n_{i}\right)\simeq 0$, yet without constraining the activation output space, enabling larger negative outputs (e.g., consider a layer where $\mu_{y}\ll 0$). For details of the trade-off, see Appendix E.
Relationship between Normalization
LayerAct functions have unique benefits that ElementAct functions cannot achieve. Therefore, it is important to select normalization methods that preserve the diversity of statistics across samples to maximize the advantages of LayerAct functions. With this consideration, batch-direction normalizations such as BN or Decorrelated Batch Normalization (DBN; Huang et al. (2018)) are identified as the most effective choices for CNNs with LayerAct functions. Conversely, other normalization methods like LN or Switch Normalization (SwitchNorm; Luo et al. (2019)) are less favored. For the details and experiments, see Appendices G, H, and I.
Experiments
In this section, we present the experimental analysis and results of LayerAct. For the detailed experimental setting to ensure reproducibility, see Appendix F.
Experimental Analysis on MNIST
In this subsection, we empirically analyze the properties of LayerAct discussed in the previous section: (1) diverse last layer representation between samples of different labels (2) activation robustness, and (3) zero-like activation mean. We trained a network with a single layer that containing 512 elements on the MNIST training dataset without any noise to observe the behavior of LayerAct functions during training.
Ensuring Diverse Representation
One of the key properties of LayerAct is that it does not produce homogeneous outputs, unlike LN, ensuring diverse representations in networks. Figure 3 shows the final layer representation of ResNet20 with LA-SiLU without normalization or BN. LayerAct has more diverse representations than ResNet20 with LN and ReLU (Figure 1), despite utilizing layer-direction normalization. This is due to the activation output, which is the product of the activation input $y$ and the activation scale $s\left(n\right)$ (see Equation 12). This mechanism allows networks with BN and LayerAct to achieve better performance on clean datasets compared to those using LN.
Figure 3: Plots of the last layer representation in ResNet20 with LA-SiLU, with no normalization (left) and BN (right).
Figure 4: Distribution of activation output fluctuation.
Activation Robustness
To confirm the robustness of LayerAct functions, we computed the activation fluctuation as defined in Definitions 3 and 5 using the network trained on the clean MNIST dataset. For the noisy input $\hat{y}_{i}$, we used two different noises with a normal distribution: one with a mean of zero and a variance of $0.5^{2}$, and another with a mean of one and a variance of $0.5^{2}$.
Figure 4 shows the distribution of the activation fluctuation with a noise distribution. Although the fluctuation distribution of the activation input was similar (see Appendix K), LayerAct functions have a significantly smaller mean and variance of activation fluctuation among the samples than any other ElementAct function in all cases. The decrease in variance is remarkable, showing that LayerAct functions are robust for all samples. Moreover, the ElementAct functions that ensure a zero-like activation mean with one-sided saturation such as SiLU or HardSiLU showed slightly larger activation fluctuations than those of ReLU or LReLU when the noise had a large mean. However, LayerAct functions maintained lower fluctuations in both cases.
Figure 5: Distribution of the activation output means of the elements in a trained network on MNIST at epoch $40$.
Zero-like Activation Mean
Figure 5 demonstrates the distribution of the activation output means of the single-layer network trained on the MNIST dataset at the epoch $40$. The distributions did not change after the epoch $40$. LayerAct functions maintain a zero-like activation mean for all epochs. Our experimental results indicate that LayerAct allows similar (before epoch $20$) or larger (after epoch $40$) negative outputs compared to other ElementAct functions. Thus, LA-SiLU and LA-HardSiLU can achieve a more zero-like activation mean than other activation functions.
CIFAR10 CIFAR10-C
Activation Clean Total Noise Blur Digital Weather Extra
ReLU 91.3 (0.045) 68.8 (1.735) 50.3 (5.967) 65.2 (2.294) 72.7 (0.948) 78.6 (0.523) 72.6 (1.471)
LReLU 91.3 (0.069) 68.7 (1.595) 49.9 (5.175) 65.3 (2.234) 72.7 (0.801) 78.4 (0.481) 72.5 (1.464)
PReLU 90.8 (0.054) 67.9 (2.010) 50.0 (6.180) 64.0 (1.990) 71.9 (1.700) 77.3 (0.669) 71.7 (1.787)
SiLU 91.5 (0.042) 69.0 (1.225) 50.2 (3.131) 65.6 (1.815) 72.5 (0.763) 78.9 (0.492) 73.0 (1.015)
HardSiLU 91.1 (0.052) 68.5 (1.453) 49.7 (5.605) 65.2 (1.971) 72.0 (0.652) 78.2 (0.418) 72.8 (1.191)
MISH 91.5 (0.053) 69.0 (1.324) 50.0 (4.568) 65.7 (1.720) 72.6 (0.627) 78.9 (0.401) 72.9 (1.267)
GELU 91.5 (0.035) 68.6 (1.290) 49.8 (3.318) 64.9 (2.311) 72.5 (0.729) 78.7 (0.276) 72.6 (1.276)
ELU 91.0 (0.030) 68.7 (1.363) 48.3 (5.193) 66.5 (1.842) 72.3 (0.605) 78.8 (0.355) 72.5 (1.220)
LA-SiLU 91.6 (0.027) 70.4 (1.392) 51.8 (3.527) 67.5 (2.155) 74.0 (0.948) 79.9 (0.559) 74.4 (1.014)
LA-HardSiLU 91.2 (0.042) 70.3 (1.384) 52.2 (5.707) 67.5 (1.238) 73.5 (0.836) 79.5 (0.501) 74.3 (1.127)
Table 2: Classification performance of ResNet20 on the CIFAR10 and CIFAR10-C. We present mean and (variance).
ImageNet ImageNet-C
Model Activation Clean Total Noise Blur Digital Weather Extra
ResNet50 ReLU 77.71 43.75 34.40 36.85 48.37 47.18 49.60
ResNet50 LReLU 77.83 43.24 32.87 36.33 48.00 47.03 49.38
ResNet50 PReLU 74.99 36.77 23.28 32.27 43.29 39.56 42.05
ResNet50 SiLU 77.85 42.31 29.74 35.94 46.59 47.36 48.77
ResNet50 HardSiLU 76.30 40.56 26.21 35.14 46.01 45.23 46.63
ResNet50 Mish 77.41 42.57 31.24 36.60 46.73 46.83 48.64
ResNet50 GELU 78.01 40.71 27.82 35.35 45.34 44.52 47.28
ResNet50 LA-SiLU 78.62 45.29 36.16 37.66 50.31 48.33 51.71
ResNet50 LA-HardSiLU 78.24 43.63 32.21 37.57 47.69 47.88 49.93
ResNet101 ReLU 79.24 46.7 36.29 39.75 52.16 49.93 52.78
ResNet101 LA-SiLU 79.12 47.87 39.62 39.85 52.82 50.84 54.16
Table 3: Classification performance of ResNet50 and ResNet101 on the clean and noisy ImageNet.
Classification Performance
We demonstrate the classification performance of LayerAct on CIFAR10, ImageNet (Russakovsky et al. 2015), and a medical image dataset. We used ResNet (He et al. 2016) as the network for our experiments on the three benchmark datasets. In our experiments on CIFAR10 and ImageNet, we utilized BN. We compared LayerAct with ReLU, LReLU, PReLU, SiLU, HardSiLU, Mish (Misra 2020), GELU (Hendrycks and Gimpel 2023), and ELU (Clevert, Unterthiner, and Hochreiter 2016). In the tables, The best results are underlined and bolded, and the second best are bolded. In the main manuscript, we present the results of ResNet20 on CIFAR10 and ResNet50 and ResNet101 on ImageNet. For more experimental results, including on CIFAR100, see Appendix J.
On CIFAR10 and CIFAR10-C
The column named clean of Table 2 presents the classification performance of ResNet20 with both LayerAct functions and ElementAct functions, benchmarked on the clean CIFAR10. We report the mean accuracy over 30 runs. LA-SiLU achieved the best performance among the activation functions. In other experiments, with ResNet32 and ResNet44 on clean CIFAR10 (Appendix J), networks with GELU achieved better results in two cases: ResNet32 on CIFAR10 and ResNet44 on CIFAR100. In the remaining combinations of networks and datasets, LA-SiLU outperformed in a significant majority of cases, especially in 39 out of 48 cases according to the $p$-value from a T-test or a Wilcoxon signed-rank test, indicating statistical significance.
To verify the robustness of LayerAct functions, we evaluated their classification performance on CIFAR10-C. The columns other except clean in Table 2 demonstrates the experimental results. The networks with LayerAct functions achieved remarkable performance compared to those with ElementAct functions. Specifically, LA-SiLU and LA-HardSiLU statistically outperformed other ElementAct functions in 46 out of 48 cases for each, according to the $p$-value from a T-test or a Wilcoxon signed-rank test. This shows that LayerAct functions exhibit greater robustness to intense noise compared to other functions. Furthermore, the experimental results on ResNets without a normalization method reveal that LA-SiLU can maintain its robustness when inputs are not excessively large.
On ImageNet
Table 3 shows the performance of ResNet50 and ResNet101 with LayerAct functions and other ElementAct functions for comparison with clean and noisy ImageNet datasets. For the column named ”clean,” we report the accuracy of 10-crop testing on the validation dataset. We used the out-of-distribution benchmark dataset, ImageNet-C (Hendrycks and Dietterich 2019), which has the same corruptions as CIFAR-C datasets. The networks with LayerAct functions outperformed those with other activation functions on all datasets. The ResNet50 with LayerAct functions, even with LA-HardSiLU which showed worse performance on the clean CIFAR10 datasets compared to SiLU or LReLU, outperformed other activation functions on clean ImageNet. Based on these results, we trained deeper networks, ResNet101, utilizing ReLU and LA-SiLU as activations. LA-SiLU maintained its superior robustness.
Activation U-net Unet++ w/o DSV Unet++ w DSV
ReLU 84.71 84.94 84.92
SiLU 84.87 85.15 85.01
LA-SiLU 85.13 85.27 85.05
Model ReLU SiLU LA-SiLU
Nat Adv Nat Adv Nat Adv
RN18 81.79 82.73 83.28 82.93 83.64 84.14
Table 4: Additional experiments on medical image segmentation (upper) and adversarial robustness (lower).
Additional Experiments
To evaluate LayerAct’s effectiveness, we conducted additional experiments on a nuclei image dataset (Goodman et al. 2018) for segmentation and on CIFAR10 to verify adversarial robustness. Table 4 shows that networks with LA-SiLU outperform those with ReLU or SiLU in both tasks. Results from segmentation tasks with U-Net (Olaf Ronneberger 2015) and UNet++ (Zhou et al. 2018) show LayerAct’s potential across different architectures for image tasks. The experimental results on adversarial robustness with ResNet18 (RN18) further underscore LayerAct’s robustness. For the detail, see Appendices L and M for experiments on medical image and adversarial robustness, respectively.
Conclusion
In this work, we identified the fundamental limitation of ElementAct in robustness and demonstrated that layer-direction normalization has the potential to enhance the robustness of activation. Inspired by our investigation, we have developed LayerAct mechanism and functions that not only enhance robustness compared to existing activation functions but also resolve the trade-off problem. Moreover, unlike LN, LayerAct does not reduce the performance of CNNs when utilized with BN on clean dataset. Experimental results on benchmark datasets show that LayerAct functions preserve essential activation properties and provide robust performance on both clean and noisy datasets, underscoring their utility and effectiveness in advanced CNN architectures.
Acknowledgements
This research was partly supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF), funded by the Ministry of Education (RS-2024-00463188) and the Ministry of Science and ICT (RS-2024-00458720), as well as by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korean government (MSIT) (No.RS-2024-00439932, SW Starlab; No.RS-2020-II201336, Artificial Intelligence Graduate School Program - UNIST; No.RS-2021-II212068, Artificial Intelligence Innovation Hub; No.RS-2024-00443780, Development of Foundation Models for Bioelectrical Signal Data and Validation of Their Clinical Applications: A Noise-and-Variability Robust, Generalizable Self-Supervised Learning Approach).
Appendix LayerAct: Advanced Activation Mechanism for Robust Inference of CNNs
Appendix A A. Important Properties of Activation
One-side saturation and zero-like activation mean stand out as crucial properties of activation for effective and efficient network training. Activation functions that exhibit one-side saturation, such as the ReLU, are favored within deep networks for their potential to provide more informative signal propagation during the backward pass by allowing for larger derivatives (Glorot, Bordes, and Bengio 2011). Furthermore, the presence of a saturation state contributes to robustness against input shifts, maintaining consistent activation outputs for elements in this state and thereby stabilizing network training (Clevert, Unterthiner, and Hochreiter 2016; Qiu, Xu, and Cai 2018).
Achieving a zero-like activation mean is another important property of activation functions, facilitating efficient network training. The activation mean refers to the average activation output of an individual element across samples. Functions such as the Exponential Linear Unit (ELU) (Clevert, Unterthiner, and Hochreiter 2016) and Parametric ReLU (PReLU) (He et al. 2015) push the activation mean toward zero by allowing negative outputs. Activation functions with this property can correct the bias shift that occurs between layers and lead to efficient network training (Clevert, Unterthiner, and Hochreiter 2016; Qiu, Xu, and Cai 2018).
The activation output of the $i^{th}$ unit of $m^{th}$ sample ($m\in\{1,2,...,M\}$) is defined as $a_{m,i}=f\left(y_{m,i}\right)$, where $f$, $y_{m,i}$, and $M$ are activation function, the $i^{th}$ activation input of the $m^{th}$ sample, and the number of samples in the batch, respectively. Ideally, a “zero-like activation mean” occurs when the activation mean of a single unit, $a_{i}$, approximates zero across the samples. Mathematically, this can be represented as:
$\begin{split}\frac{1}{M}\sum^{M}_{m=1}a_{m,i}\approx 0\end{split}.$
However, approximating the activation mean to zero is challenging for activation functions that saturate at (large) negative outputs, such as ELU, SiLU, or FReLU. Due to this saturation, previous studies have defined the “zero-like activation mean” property of an activation function as its ability to “push” the activation mean towards zero. In mathematical terms, this can be presented as $\left|\mu_{a_{i}}\right|\ll c$, where $c$ is a small positive constant (Clevert, Unterthiner, and Hochreiter 2016; Qiu, Xu, and Cai 2018).
Appendix B B. Detailed Derivation Process of Activation Fluctuation with LN
In this section, we present the detailed derivation process of the equations in Definition 4 in the main manuscript as follows:
$\left\|f\left(\hat{n}\right)-f\left(n\right)\right\|=$ $\sum^{D}_{i=1}\left|\hat{n}_{i}s\left(\hat{n}_{i}\right)-n_{i}s\left(n_{i}\right)\right|$
$=$ $\sum^{D}_{i=1}\left|\frac{\hat{y}_{i}+\epsilon_{i}-\mu_{y}-\mu_{\epsilon}}{\sqrt{\sigma^{2}_{y}+\sigma^{2}_{\epsilon}+\alpha}}s\left(\hat{n}_{i}\right)-\frac{y_{i}-\mu_{y}}{\sqrt{\sigma^{2}_{y}+\alpha}}s\left(n_{i}\right)\right|$
$\approx$ $\sum^{D}_{i=1}\left|n_{i}\left(s\left(\hat{n}_{i}\right)-s\left(n_{i}\right)\right)+\frac{\left(\epsilon-\mu_{\epsilon}\right)s\left(\hat{n}_{i}\right)}{\sqrt{\sigma_{y}^{2}+\alpha}}\right|$
$\leq$ $\sum^{D}_{i=1}\left(\left|n_{i}\right|\cdot\left|s\left(\hat{n}_{i}\right)-s\left(n_{i}\right)\right|+\frac{\left|\epsilon-\mu_{\epsilon}\right|\cdot s\left(\hat{n}_{i}\right)}{\sqrt{\sigma_{y}^{2}+\alpha}}\right),$
where $\sqrt{\sigma_{y}^{2}+\alpha+\sigma_{\epsilon}^{2}}\approx\sqrt{\sigma_{y}^{2}+\alpha}>1$ when $\sigma_{y}\gg\sigma_{\epsilon}$ and $\alpha$ is sufficiently large.
Appendix C C. Difference between LayerAct and Activation with LN
In this section, we provide detailed comparison between the activation outputs of LayerAct and those of element-level activation functions paired with LN. When a LN layer is placed right before an activation layer, the $i^{th}$ activation output is $a_{i}=n^{LN}_{i}s\left(n^{LN}_{i}\right)$, where $n^{LN}_{i}$ is the $i^{th}$ normalized output of LN. Conversely, the $i^{th}$ activation output of a LayerAct function is $a_{i}=y_{i}s\left(n_{i}\right)$, as defined in Equation 12 in the main manuscript.
The critical distinction between activation with LN and LayerAct lies in the preservation of input mean and variance statistical information in the activation output. LayerAct can preserve the statistical information, but activation with LN cannot. With LN, the activation function takes a normalized input in layer-direction, resulting in activation outputs that exhibit similar mean and variance across samples (as shown in the activation output equation for LN above). This homogenization of statistical information across samples, a characteristic of LN, is a reason why BN often outperforms LN in non-sequential models such as CNNs (Labatie et al. 2021; Lubana, Dick, and Tanaka 2021). In the main manuscript, we demonstrated that this homogeneity limitation of LN leads the representation of the samples with different labels to be similar (see Figure 1 in the main manuscript).
LayerAct, on the other hand, produces more distinguishable activation outputs between samples by preserving statistical variation between samples. This is due to the fact that only the activation scale function of LayerAct uses the layer-normalized input, not the LayerAct function itself (as shown in Equation 12 in the main manuscript). Based on this unique output of LayerAct, networks with LayerAct can produce much more distinguishable representation of samples compared to those with activation and LN (see Figure 3 in the main manuscript).
Figure 6: LA-SiLU with different mean and variance values in the input. The distribution of the activation input is: (1) $\mu_{y}=0$, $\sigma_{y}=1$, (2) $\mu_{y}=0$, $\sigma_{y}=5$, (3) $\mu_{y}=-5$, $\sigma_{y}=1$, and (4) $\mu_{y}=5$, $\sigma_{y}=1$ from the left to right.
Figure 7: LA-HardSiLU with different mean and variance values in the input. The distribution of the activation input is: (1) $\mu_{y}=0$, $\sigma_{y}=1$, (2) $\mu_{y}=0$, $\sigma_{y}=5$, (3) $\mu_{y}=-5$, $\sigma_{y}=1$, and (4) $\mu_{y}=5$, $\sigma_{y}=1$ from the left to right.
Appendix D D. Activation Output of LayerAct Functions
In this section, we provide detailed comparison between the activation outputs of LayerAct and those of element-level activation functions paired with LN. When a LN layer is placed right before an activation layer, the $i^{th}$ activation output is $a_{i}=n^{LN}_{i}s\left(n^{LN}_{i}\right)$, where $n^{LN}_{i}$ is the $i^{th}$ normalized output of LN. Conversely, the $i^{th}$ activation output of a LayerAct function is $a_{i}=y_{i}s\left(n_{i}\right)$, as defined in Equation 12 in the main manuscript.
The critical distinction between activation with LN and LayerAct lies in the preservation of input mean and variance statistical information in the activation output. LayerAct can preserve the statistical information, but activation with LN cannot. With LN, the activation function takes a normalized input in layer-direction, resulting in activation outputs that exhibit similar mean and variance across samples (as shown in the activation output equation for LN above). This homogenization of statistical information across samples, a characteristic of LN, is a reason why BN often outperforms LN in non-sequential models such as CNNs (Labatie et al. 2021; Lubana, Dick, and Tanaka 2021).
LayerAct, on the other hand, produces more distinguishable activation outputs between samples by preserving statistical variation between samples. This is due to the fact that only the activation scale function of LayerAct uses the layer-normalized input, not the LayerAct function itself (as shown in Equation 12 in the main manuscript).
Appendix E E. Trade-off between Two Properties
A fundamental trade-off exists between the important properties of activation, one-side saturation and zero-like mean activation. Activation functions that have a saturation state typically limit negative outputs, consequently driving the activation mean away from zero. On the other hand, to achieve a zero-like activation mean, certain functions like PReLU allow large negative outputs. However, functions with this allowance are not robust due to the absence of saturation.
To address this trade-off issue, various activation functions, such as Sigmoid Linear Unit (SiLU, also known as SWISH) (Elfwing, Uchibe, and Doya 2018; Ramachandran, Zoph, and Le 2018), Gaussian Error Linear Unit (GELU) (Hendrycks and Gimpel 2023), and Mish (Misra 2020), saturate the large negative outputs only. These activation functions can achieve a zero-like activation mean with small negative outputs. However, a trade-off still exists because the restriction of negative outputs, designed to ensure saturation, prevents the allowance of large negative outputs, thereby restraining the mean of activation from being more zero-like.
Additionally, this trade-off cannot be addressed by LN. Even when a LN layer is placed before the activation layer, the large negative output is restricted with element-level activation. Therefore, mitigating the trade-off between activation properties is a distinct benefit of LayerAct functions that element-level activation functions cannot attain.
Appendix F F. Experimental Setting
We implemented LayerAct functions and networks for experiment with PyTorch (Paszke et al. 2019). All networks used in our experiments were trained on NVIDIA A100. We used multiple devices to train the networks on ImageNet, and a single device for the other experiments. The versions of Python and the packages were (1) Python 3.9.12, (2) numpy 1.19.5 (3) PyTorch 1.11.0, and (4) torchvision 0.12.0. We used cross entropy loss functions for all the experiments. The random seeds of the experiments were $11\times i$ where $i\in\{1,2,...,30\}$ on CIFAR10 and CIFAR100 and $11$ on ImageNet.
To train networks on MNIST for experimental analysis, we applied batch gradient descent for $80$ epochs with the weight decay and momentum fixed to $0.0001$ and $0.9$, respectively. The learning rate started from 0.01, and was multiplied $0.1$ at epochs $40$ and $60$ as scheduled. We used ResNet (He et al. 2016) with BN right before activation for experiments on CIFAR10, CIFAR100 and ImageNet. We initialized the weights following the methods proposed by He et al. (He et al. 2015). For all experiments, the weight decay, momentum, and initial learning rate were $0.0001$, $0.9$ and $0.1$, respectively.
For CIFAR10 and CIFAR100, we trained ResNet20, ResNet32, and ResNet44 with a basic block using the stochastic gradient descent with a batch size of 128 for about 64000 iterations. We randomly selected $10\%$ of the training dataset as the validation set. The learning rate was scheduled to decrease by the factor of 10 at 32000 and 48000 iterations. For the data augmentation of CIFAR10 and CIFAR100, we followed (Lee et al. 2015). We rescaled the data between 00 and $1$, padded $4$ pixels on each side, and randomly sampled a $32\times 32$ crop from the padded image or its horizontal flip. The data was normalized after augmentation. For testing, we did not apply data augmentation, only normalized the data. The hyper-parameter $\alpha$ of LayerAct functions for the experiments was set to $0.00001$.
For the experiment with ImageNet, we trained ResNet50 and ResNet101 with the bottleneck block using stochastic gradient descent, and the batch size was 256 for about 600000 iterations. The learning rate was scheduled to decrease by a factor of 10 at 180000, 360000, and 540000 iterations. For the data augmentation on ImageNet, we rescaled the data between 00 and $1$, resized it to $224\times 244$, and randomly sampled a $224\times 224$ crop from an image or its horizontal flip (Krizhevsky, Sutskever, and Hinton 2012). We normalized the data after data augmentation. For testing, we resized the data to be $256\times 256$ and applied 10-crop. Afterward, the data was normalized. To ensure stable learning, we set the hyper-parameter $\alpha$ of LayerAct functions to $0.1$ which is larger than those for CIFAR10 and CIFAR100.
For the reproducibility, we have attached the code that used for the experiments in supplementary materials.
Appendix G G. Relationship between LayerAct and BN
In our main manuscript, we emphasized the importance of selecting an appropriate normalization method for employing LayerAct functions effectively. This section delves into the relationship between LayerAct and BN, the most dominant normalization method for CNNs. In this section, our objective is to discuss the benefits of BN in CNNs and demonstrate how LayerAct functions collaborate well with BN, in contrast to LN.
BN operates along three dimensions, batch, height, and width, particularly for image datasets, where the input matrix $y\in\mathbb{R}^{B\times C\times H\times W}$. To simplify, we consider a flattened dimension that combines height and width, denoted by size $D$. The mean, variance, and output of BN are then defined as follows:
$\mu^{BN}_{j}=$ $\frac{1}{BD}\sum^{B}_{i=1}\sum^{D}_{k=1}y_{i,j,k},$ (17)
$\sigma^{BN}_{j}=$ $\frac{1}{BD}\sum^{B}_{i=1}\sum^{D}_{k=1}\left(y_{i,j,k}-\mu^{BN}_{j}\right),$ (18)
$n^{BN}_{i,j,k}=$ $\gamma_{j}\cdot\frac{y_{i,j,k}-\mu^{BN}_{j}}{\sqrt{{\sigma^{BN}_{j}}^{2}+\alpha}}+\beta_{j}.$ (19)
BN enjoys the common advantages of normalization methods, re-scaling and re-centering operations that significantly enhance the efficiency and effectiveness of network training, promote stable training, and allow the use of a larger learning rate (Lubana, Dick, and Tanaka 2021; Labatie et al. 2021; Ioffe and Szegedy 2015).
Additionally, BN offers the unique benefit of avoiding channel collapse, a phenomenon where a channel exhibits a linear activation (i.e. a channel loses its non-linear activation), leading to inefficient network training (Labatie et al. 2021). For instance, when a channel’s activation inputs are consistently positive or negative across all samples, the activation output of a ReLU layer becomes linear, $a=y$ or $a=0$, respectively. BN addresses this by normalizing across the batch dimension, ensuring a diverse distribution of channel activations among samples. Moreover, it does not have the homogenization issue present in LN.
When utilized with LayerAct, BN does not compromise the benefits of LayerAct because they operate on different normalization dimensions. Given that the output of LayerAct functions is the product of the activation input (i.e. the normalization output) and the activation scale, BN’s benefits extend to the activation output, enhancing overall network performance as when utilized with element-level activation function. Therefore, utilizing LayerAct with BN preserves the advantages of both, enhancing the efficiency of network training and robustness of network inference.
In summary, the choice of normalization method for CNNs with LayerAct functions should be made carefully, taking into account the interplay between LayerAct and the normalization. While LN may diminish LayerAct’s benefits, as discussed in Appendix H, BN, with its beneficial properties, would be a suitable choice for CNNs with LayerAct functions on image datasets.
Table 5: Classification performance on CIFAR and CIFAR-C datasets without a normalization method. C10 and C100 denotes CIFAR10 and CIFAR100, respectively.
CIFAR CIFAR-C
Data Model Activation Clean Total Noise Blur Digital Weather Extra
C10 ResNet20 ReLU 88.51 67.43 53.27 62.12 70.93 75.91 71.38
C10 ResNet20 SiLU 88.29 68.94 54.66 65.84 71.6 76.54 72.48
C10 ResNet20 LA-SiLU 88.95 69.36 54.71 65.58 71.93 77.59 73.34
C10 ResNet32 ReLU 89.56 68.94 54.48 63.90 72.45 77.52 72.73
C10 ResNet32 SiLU 46.51 38.69 33.21 37.52 39.62 41.6 40.11
C10 ResNet32 LA-SiLU 89.12 70.59 56.29 66.81 73.11 78.94 74.22
C10 ResNet44 ReLU 90.03 69.97 55.75 65.08 73.38 78.45 73.62
C10 ResNet44 SiLU 15.25 14.18 13.5 14.01 14.26 14.55 14.39
C10 ResNet44 LA-SiLU 88.77 71.65 58.15 68.32 74.01 79.49 74.88
C100 ResNet20 ReLU 59.46 37.69 21.17 38.76 41.36 44.15 38.89
C100 ResNet20 SiLU 59.74 40.39 24.21 42.72 43.51 46.45 41.01
C100 ResNet20 LA-SiLU 61.01 42.44 26.07 44.50 45.74 48.8 42.98
C100 ResNet32 ReLU 60.71 39.56 23.05 40.78 43.23 45.93 40.67
C100 ResNet32 SiLU 20.36 14.99 10.05 16.27 15.89 16.47 15.02
C100 ResNet32 LA-SiLU 60.08 44.01 28.77 46.73 47.02 49.51 44.19
C100 ResNet44 ReLU 61.59 40.88 24.57 42.08 44.60 47.16 41.90
C100 ResNet44 SiLU 2.91 2.43 1.97 2.58 2.53 2.54 2.42
C100 ResNet44 LA-SiLU 60.30 44.26 29.22 46.74 47.45 49.65 44.50
Table 6: Classification performance on CIFAR and CIFAR-C datasets with LN. C10 and C100 denote CIFAR10 and CIFAR100, respectively.
CIFAR CIFAR-C
Data Model Activation Clean Total Noise Blur Digital Weather Extra
C10 ResNet20 ReLU 88.24 73.77 62.05 71.26 76.48 79.69 76.46
C10 ResNet20 SiLU 87.53 74.3 63.32 71.74 77.33 79.77 76.58
C10 ResNet20 LA-SiLU 88.52 75.31 63.81 73.11 78.23 80.92 77.6
C10 ResNet32 ReLU 88.55 74.48 63.33 71.92 76.94 80.33 77.11
C10 ResNet32 SiLU 87.27 75.3 64.78 72.92 78.55 80.3 77.32
C10 ResNet32 LA-SiLU 87.85 76.39 67.35 74.22 78.80 80.99 78.32
C10 ResNet44 ReLU 88.58 75.2 64.32 72.75 77.87 80.67 77.67
C10 ResNet44 SiLU 86.65 75.11 65.67 72.73 77.84 79.85 77.1
C10 ResNet44 LA-SiLU 86.88 76.73 69.53 74.98 78.51 80.43 78.43
C100 ResNet20 ReLU 61.30 44.04 31.82 44.39 47.03 48.86 45.07
C100 ResNet20 SiLU 60.45 45.93 35.66 46.46 48.46 49.91 46.59
C100 ResNet20 LA-SiLU 62.30 45.30 31.63 45.90 48.72 50.65 46.17
C100 ResNet32 ReLU 63.21 45.92 33.23 46.17 49.26 50.94 46.84
C100 ResNet32 SiLU 60.04 47.14 37.37 48.60 49.57 50.33 47.41
C100 ResNet32 LA-SiLU 60.76 47.26 36.63 48.51 49.97 51.03 47.51
C100 ResNet44 ReLU 64.18 46.63 33.92 46.89 49.68 51.72 47.75
C100 ResNet44 SiLU 59.58 47.41 38.28 48.69 49.45 50.61 47.73
C100 ResNet44 LA-SiLU 60.17 47.31 38.10 48.82 49.63 50.26 47.46
Appendix H H. LayerAct with no normalization or LN
In this section, we explore the cases when LA-SiLU is utilized with LN. For the further investigation, we first examine the relationship between LN and LayerAct, subsequently presenting experimental results on CIFAR datasets for two scenarios: (1) networks without any normalization methods, and (2) networks with LN. ReLU and SiLU were selected as baseline element-level activation functions, with ReLU being the most popular choice for CNNs, and SiLU being as the corresponding element-level activation function of LA-SiLU.
For the networks incoporating LN, the normalization output of a LN layer is expressed as $n^{\hat{LN}}=\gamma\cdot\frac{y-\mu_{y}^{2}}{\sqrt{\sigma_{y}^{2}+\alpha}}+\beta$. This equation results in the output of SiLU $n^{\hat{LN}}s\left(n^{\hat{LN}}\right)$ and LA-SiLU $n^{\hat{LN}}s\left(n\right)$ becoming similar. When the learnable parameter $\gamma$ and $\beta$ are $1$ and 00, respectively, the outputs $n^{\hat{LN}}s\left(n^{\hat{LN}}\right)$ and $n^{\hat{LN}}s\left(n\right)$ become identical, nullifying any distinct advantage offered by the LayerAct mechanism. Thus, the more the output of the LN layer approximates the normalized input of LayerAct’s activation scale function (i.e., the more gain and bias approximate to $1$ and 00, respectively), the more the benefits of LayerAct are reduced.
Nevertheless, the benefits of LayerAct functions, addressing the trade-off between two important activation properties and potentially exhibiting lower variance in robustness across samples, can still be partially retained when combined with LN that the learnable parameters $\gamma$ and $\beta$ are far from $1$ and 00, respectively. This is because the layer-direction normalized input of LayerAct’s activation scale function does not utilize an affine function. By not utilizing an affine function, we can ensure that $\left\|s\left(\hat{n}\right)\right\|$ of LayerAct functions remains small enough for robustness. When an affine function is used, the mean of the normalized vector $n$ can become large if the parameter $\beta$ is large. Such large $n$ leads to a large $\left\|s\left(\hat{n}\right)\right\|$, which does not ensure robustness.
Additionally, LayerAct functions can allow a larger negative output compared to element-level activation functions, as they do not restrict large negative outputs. Furthermore, the robustness benefit of LayerAct becomes significant when the normalized outputs of LN (i.e. activation input) are mostly outside the saturation state.
To investigate this, we conducted experiments on ResNets both without normalization and with LN. We followed the same experimental setting of the experiments in our main manuscript. The results reported in the tables are the average accuracies over 30 runs, each with different seeds for weight initialization.
Table 5 presents the performance of ResNets without any normalization methods on both clean and noisy CIFAR datasets, with a learning rate set at 0.01 for stable training. The results demonstrate that networks with LA-SiLU outperformed those with ReLU and SiLU on noisy datasets. This indicates that the robustness of LayerAct functions is independent from the presence of normalization layers. Additionally, while networks with SiLU, the corresponding element-level activation function, experienced training instability and often exploded, those with LA-SiLU maintained stable training. This suggests that the LayerAct mechanism can contribute to more stable network training. However, in deeper networks such as ResNet32 and ResNet44, the performance with ReLU surpassed those with LA-SiLU. This outcome may be because of the more complex representation of LA-SiLU’s output, pointing towards the necessity of a normalization layer to mitigate potential overfitting issues.
Table 6 demonstrates the performance of ResNets with LN on both clean and noisy CIFAR10 and CIFAR100 datasets. It validated our concerns about the relationship between LN and LayerAct. On noisy CIFAR100, ResNet20 and ResNet44 with SiLU demonstrated similar or greater robustness than those with LA-SiLU. This suggests that the robustness benefits of LayerAct are diminished when utilized with LN. However, it is important to note that the benefits of LayerAct may still be partially preserved even when LN is applied. For clean and noisy CIFAR10, the results where similar with those in Table 5, the networks with LA-SiLU demonstrated more robust inference compared to those with other activation functions on noisy datasets, while the performance of the deeper networks (ResNet32 and ResNet44) with ReLU where better on clean datasets than those with LA-SiLU. Furthermore, on clean datasets, the networks with LA-SiLU outperformed those with SiLU, which employs the same function (Logistic Sigmoid) for the activation scale with LA-SiLU. This indicates that the advantage of LA-SiLU, in addressing the trade-off, can enhance network performance.
Additionally, networks employing LayerAct functions with BN, which normalizes in a different dimension from that of LayerAct functions and remains the dominant normalization method for CNNs, perform effectively. In our experiments reported in the main manuscript, BN was employed for all networks, except for UNets, which do not utilize any normalization layer. The networks with LayerAct functions exhibited remarkable performance on both clean and noisy datasets. Consequently, despite the necessity for careful selection of the normalization method when using LayerAct functions, their robustness and improved performance when combined with BN make them an attractive choice for CNNs.
Table 7: Classification performance on CIFAR and CIFAR-C datasets with SN. We report the average mean accuracy in the table. C10 and C100 denotes CIFAR10 and CIFAR100, respectively.
CIFAR CIFAR-C
Data Model Activation Clean Total Noise Blur Digital Weather Extra
C10 ResNet20 ReLU 89.65 72.40 56.07 71.35 75.24 80.43 74.81
C10 ResNet20 SiLU 90.60 74.17 57.70 72.94 77.57 82.21 76.33
C10 ResNet20 LA-SiLU 89.56 72.43 55.9 71.05 75.47 80.87 74.73
C10 ResNet32 ReLU 90.70 73.90 58.24 72.84 76.3 81.79 76.42
C10 ResNet32 SiLU 90.79 74.65 58.78 73.39 77.64 82.57 76.91
C10 ResNet32 LA-SiLU 89.94 72.72 56.15 71.07 75.52 81.48 75.23
C10 ResNet44 ReLU 91.40 74.65 58.10 74.01 77.18 82.81 76.99
C10 ResNet44 SiLU 74.48 61.70 49.50 60.58 64.11 67.87 63.41
C10 ResNet44 LA-SiLU 89.36 72.02 54.85 70.59 75.07 81.05 74.24
C100 ResNet20 ReLU 57.36 37.01 20.61 38.43 39.75 43.56 38.59
C100 ResNet20 SiLU 64.55 42.96 25.13 43.91 46.35 50.42 44.55
C100 ResNet20 LA-SiLU 63.88 41.76 23.23 42.83 45.28 49.43 43.4
C100 ResNet32 ReLU 60.19 38.81 21.89 40.02 41.30 46.00 40.61
C100 ResNet32 SiLU 64.35 42.45 25.00 43.08 45.86 49.92 44.00
C100 ResNet32 LA-SiLU 64.05 42.27 23.72 43.46 45.50 50.16 43.89
C100 ResNet44 ReLU 61.64 39.93 22.64 41.03 42.47 47.54 41.67
C100 ResNet44 SiLU 44.8 29.57 18.11 29.96 31.98 34.21 30.73
C100 ResNet44 LA-SiLU 62.99 41.91 22.86 43.06 46.02 49.79 43.06
Appendix I I. LayerAct with Other Normalization Methods.
To investigate the relationship between LayerAct and various normalization methods, we conducted experiments on ResNets with Switchable Normalization (SN) (Luo et al. 2019), Instance enhancement batch normalization (IEBN) (Liang et al. 2020), and Decorrelated Batch Normalization (DBN) (Huang et al. 2018). We used the same experiment setting with those of our main manuscript. We report the average accuracy over 10 runs for the experiments.
Analysis of the experimental results revealed that LayerAct functions are not effectively compatible with normalizations that can cause a large variance in the channel means, such as SN and IEBN. This incompatibility arises because LayerAct is more sensitive to such large variance between channel means, which causes channels with smaller means to be more likely to become inactivated compared to those with larger means.
Table 7 demonstrates that utilizing LayerAct with SN is not effective. This ineffectiveness arises because SN is a combination of BN, LN, and Instance normalization (IN) (Ulyanov, Vedaldi, and Lempitsky 2017), while the benefit of LayerAct functions diminish when preceded by LN. Additionally, LayerAct functions are more sensitive to the presence of similar channel characteristics across samples compared to element-level activation functions because the channel dimension is incorporated into normalizing dimension of LayerAct’s activation scale input. IN aggressively normalizes the mean and variance across channels towards uniformity. Therefore, given the composite normalization strategy of SN, we expect that the effect of IN will lead to a homogenization of channel characteristics, thus undermining the efficiency of LayerAct functions.
Table 8: Classification performance on CIFAR and CIFAR-C datasets with IEBN. We report the average mean accuracy in the table. C10 and C100 denotes CIFAR10 and CIFAR100, respectively.
CIFAR CIFAR-C
Data Model Activation Clean Total Noise Blur Digital Weather Extra
C10 ResNet20 ReLU 91.50 70.12 52.49 66.73 73.44 79.64 73.89
C10 ResNet20 SiLU 91.44 68.98 49.39 66.26 72.29 79.00 73.06
C10 ResNet20 LA-SiLU 91.63 70.64 52.70 67.30 74.20 79.87 74.65
C10 ResNet32 ReLU 92.54 71.67 54.06 68.82 74.82 81.12 75.11
C10 ResNet32 SiLU 91.85 71.07 53.39 68.26 73.90 80.50 74.90
C10 ResNet32 LA-SiLU 92.36 71.77 54.19 68.19 75.32 81.15 75.60
C10 ResNet44 ReLU 92.78 72.18 55.20 68.83 75.61 81.59 75.43
C10 ResNet44 SiLU 92.08 71.23 52.79 68.54 74.27 80.95 74.99
C10 ResNet44 LA-SiLU 92.50 73.01 56.97 69.18 76.03 82.34 76.5
C100 ResNet20 ReLU 66.62 41.97 22.81 42.51 45.07 50.30 44.37
C100 ResNet20 SiLU 66.19 40.88 21.14 41.08 44.29 49.68 43.28
C100 ResNet20 LA-SiLU 66.73 41.59 21.31 41.98 45.26 50.53 43.82
C100 ResNet32 ReLU 68.17 43.49 23.81 43.89 46.76 52.30 45.78
C100 ResNet32 SiLU 67.43 42.30 23.15 42.38 45.33 51.14 44.68
C100 ResNet32 LA-SiLU 67.97 43.56 24.49 43.39 46.94 52.47 45.76
C100 ResNet44 ReLU 69.47 44.73 25.74 44.77 47.77 53.51 47.12
C100 ResNet44 SiLU 68.18 43.47 24.76 43.37 46.46 52.22 45.89
C100 ResNet44 LA-SiLU 68.41 45.29 26.89 45.15 48.67 54.03 47.13
Table 8 demonstrates the performance of networks with IEBN. With the exception of ResNet20 on CIFAR100, networks with LA-SiLU demonstrated enhanced performance on noisy datasets when compared to their counterparts utilizing ReLU and SiLU. Conversely, ReLU outperformed LA-SiLU in ResNet32 and ResNet44 models. Nonetheless, it is important to highlight that LA-SiLU consistently surpassed SiLU, which utilizes the same activation scale function, across all tested scenarios. These results imply that the mechanism of LayerAct holds promise for enhancing efficiency. It is also important to consider careful consideration and integration of network complexity, particularly due to the interplay between normalization and activation functions, are essential, given that the scale function of ReLU is considerably simpler compared to sigmoid, the scale function of LA-SiLU and SiLU.
Table 9: Classification performance on CIFAR and CIFAR-C with DBN. We report the average mean accuracy in the table. C10 and C100 denotes CIFAR10 and CIFAR100, respectively.
CIFAR CIFAR-C
Data Model Activation Clean Total Noise Blur Digital Weather Extra
C10 ResNet20 ReLU 90.8 65.52 35.87 65.43 71.27 78.68 68.93
C10 ResNet20 SiLU 87.69 63.55 38.16 61.21 70.24 74.77 67.01
C10 ResNet20 LA-SiLU 91.72 69.36 45.68 67.81 73.41 80.6 73.37
C10 ResNet32 ReLU 91.85 68.1 37.15 68.38 75.07 81.04 71.1
C10 ResNet32 SiLU 92.22 68.3 40.01 67.24 73.98 81.44 71.77
C10 ResNet32 LA-SiLU 92.27 70.36 45.62 69.32 74.58 81.96 74.14
C10 ResNet44 ReLU 92.35 69.79 39.25 70.48 76.61 82.31 72.67
C10 ResNet44 SiLU 92.39 69.48 42.19 68.97 74.63 81.63 73.13
C10 ResNet44 LA-SiLU 92.36 70.17 44.0 69.35 74.63 82.22 74.11
C100 ResNet20 ReLU 58.38 34.19 13.48 34.26 38.28 43.67 36.07
C100 ResNet20 SiLU 57.82 32.13 13.41 31.57 35.78 40.99 34.22
C100 ResNet20 LA-SiLU 58.19 32.24 14.32 30.95 35.11 41.4 34.96
C100 ResNet32 ReLU 54.06 30.29 14.16 28.67 34.09 38.34 32.14
C100 ResNet32 SiLU 62.0 35.26 14.99 34.41 39.35 45.01 37.46
C100 ResNet32 LA-SiLU 65.02 38.75 19.18 37.84 41.81 48.66 41.36
C100 ResNet44 ReLU 60.32 34.64 14.96 33.68 39.03 44.1 36.51
C100 ResNet44 SiLU 64.67 37.58 15.81 37.63 42.01 47.2 39.83
C100 ResNet44 LA-SiLU 67.59 42.44 21.03 42.54 46.52 52.15 44.6
Table 9 demonstrates the performance of networks with DBN. Except for ResNet20 on CIFAR100, networks with LA-SiLU showed similar or improved performance on clean datasets, and outstanding performance on noisy datasets compared to those employing ReLU and SiLU. The results of these experiments highlight the potential applicability of LayerAct functions in conjunction with advanced batch-direction normalization methods.
] CIFAR10 CIFAR100 Activation ResNet20 ResNet32 ResNet44 ResNet20 ResNet32 ResNet44 ReLU 91.29 92.03 92.03 65.92 67.04 68.02 LReLU 91.31 92.03 92.03 65.88 67.37 67.96 PReLU 90.82 92.03 - 64.00 66.35 67.68 SiLU 91.45 92.17 92.18 65.89 67.22 67.71 HardSiLU 91.09 91.77 91.42 65.19 66.49 66.38 Mish 91.48 92.21 92.30 65.85 67.18 68.06 GELU 91.50 92.25 92.22 65.84 67.30 68.19 ELU 91.04 91.61 91.68 66.24 67.01 67.55 LA-SiLU 91.60 92.20 92.36 66.39 67.74 68.07 LA-HardSiLU 91.21 91.68 91.36 66.16 66.63 65.51
Table 10: Classification performance on the clean CIFAR10 and CIFAR100
Appendix J J. Additional Tables
Table 10 shows the classification performance of networks on clean CIFAR10 and CIFAR100 datasets. LA-SiLU outperformed element-level activation functions in four out of six cases. Specifically, LA-SiLU always demonstrated superior performance compared to SiLU.
] Data Model Activation Total Noise Blur Digital Weather Extra CIFAR10-C ResNet32 ReLU 72.00 53.07 67.62 74.75 80.44 74.41 CIFAR10-C ResNet32 LReLU 72.01 52.66 67.86 74.77 80.42 74.5 CIFAR10-C ResNet32 PReLU 71.7 52.82 67.06 74.72 79.90 74.17 CIFAR10-C ResNet32 SiLU 71.7 52.37 67.21 74.08 80.51 74.38 CIFAR10-C ResNet32 HardSiLU 71.32 52.75 66.75 73.39 79.66 74.28 CIFAR10-C ResNet32 Mish 71.96 53.09 67.42 74.3 80.57 74.64 CIFAR10-C ResNet32 GELU 71.64 52.44 67.18 74.11 80.26 74.26 CIFAR10-C ResNet32 LA-SiLU 72.8 54.02 68.42 75.13 81.36 75.51 CIFAR10-C ResNet32 LA-HardSiLU 72.6 55.36 67.71 74.70 80.53 75.62 CIFAR10-C ResNet44 ReLU 73.71 56.39 70.03 76.05 81.27 75.92 CIFAR10-C ResNet44 LReLU 73.69 56.03 70.10 76.09 81.43 75.81 CIFAR10-C ResNet44 SiLU 72.45 53.12 68.64 74.80 80.88 75.06 CIFAR10-C ResNet44 HardSiLU 72.63 55.51 68.72 74.73 79.86 75.34 CIFAR10-C ResNet44 Mish 72.79 53.74 68.86 75.22 81.18 75.3 CIFAR10-C ResNet44 GELU 72.82 54.65 68.69 75.26 80.85 75.24 CIFAR10-C ResNet44 LA-SiLU 73.5 55.29 69.14 75.73 81.91 76.19 CIFAR10-C ResNet44 LA-HardSiLU 73.33 57.45 68.48 75.30 80.64 76.30 CIFAR100-C ResNet32 ReLU 43.51 24.10 41.99 45.81 50.58 44.32 CIFAR100-C ResNet32 LReLU 43.58 23.8 42.12 45.94 50.73 44.42 CIFAR100-C ResNet32 PReLU 42.44 23.72 40.31 44.81 49.42 43.27 CIFAR100-C ResNet32 SiLU 42.94 23.06 41.27 45.01 50.31 44.02 CIFAR100-C ResNet32 HardSiLU 42.67 23.64 40.87 44.87 49.54 43.71 CIFAR100-C ResNet32 Mish 42.95 22.69 41.53 45.05 50.37 43.97 CIFAR100-C ResNet32 GELU 43.00 23.15 41.45 45.21 50.21 43.94 CIFAR100-C ResNet32 LA-SiLU 44.6 24.16 43.58 46.74 52.11 45.51 CIFAR100-C ResNet32 LA-HardSiLU 44.86 25.98 43.83 47.02 51.50 45.81 CIFAR100-C ResNet44 ReLU 44.77 25.45 43.05 47.33 51.94 45.44 CIFAR100-C ResNet44 LReLU 44.8 25.57 43.26 47.17 51.91 45.5 CIFAR100-C ResNet44 PReLU 44.31 25.78 42.19 46.7 51.44 44.97 CIFAR100-C ResNet44 SiLU 44.04 24.52 42.61 46.16 51.08 45.01 CIFAR100-C ResNet44 HardSiLU 44.11 26.22 42.77 46.29 50.32 44.9 CIFAR100-C ResNet44 Mish 44.14 24.18 42.69 46.36 51.37 45.11 CIFAR100-C ResNet44 GELU 43.93 24.39 42.29 46.15 51.02 44.85 CIFAR100-C ResNet44 LA-SiLU 46.12 26.68 44.94 48.32 53.44 46.89 CIFAR100-C ResNet44 LA-HardSiLU 46.83 30.85 45.83 48.93 52.5 47.35
Table 11: Classification performance of ResNet32 and ResNet44 on noisy CIFAR10 and CIFAR100. We report the average mean accuracy over 30 runs.
Table 12: Standard deviation of ResNet20s’ accuracy on CIFAR10 and CIFAR10-C with different activation functions. We average in the table.
Activation Clean Total Noise Blur Digital Weather Extra
ReLU 0.2120 1.3171 2.4428 1.5145 0.9736 0.7231 1.2127
LReLU 0.2635 1.2630 2.2750 1.4947 0.8948 0.6938 1.2099
PReLU 0.2323 1.4176 2.4859 1.4106 1.3038 0.8178 1.3369
SiLU 0.2057 1.1067 1.7695 1.3471 0.8738 0.7012 1.0076
HardSiLU 0.2287 1.2052 2.3675 1.4039 0.8074 0.6465 1.0914
MISH 0.2302 1.1507 2.1373 1.3116 0.7921 0.6335 1.1255
GELU 0.1868 1.1358 1.8216 1.5201 0.8537 0.5251 1.1298
ELU 0.1742 1.1673 2.2787 1.3572 0.7776 0.5961 1.1047
LA-SiLU 0.1645 1.1800 1.8781 1.4681 0.9738 0.7475 1.0069
LA-HardSiLU 0.2051 1.1765 2.3889 1.1128 0.9142 0.7081 1.0616
Table 13: Standard deviation of ResNet32s’ accuracy on CIFAR10 and CIFAR10-C with different activation functions. We average in the table.
Activation Clean Total Noise Blur Digital Weather Extra
ReLU 0.3880 1.5955 2.8978 1.7971 1.3270 0.9254 1.3556
LReLU 0.3356 1.7599 3.0043 2.1394 1.4202 1.0200 1.5270
PReLU 0.2809 1.6875 2.6927 1.8571 1.6496 0.9261 1.5634
SiLU 0.2188 1.3848 2.7176 1.7485 0.8207 0.7220 1.2483
HardSiLU 0.2200 1.2341 2.0973 1.5508 0.8625 0.8104 1.0653
MISH 0.2497 1.2869 2.1864 1.7262 0.8797 0.6475 1.2194
GELU 0.1568 1.2598 2.3925 1.4798 0.8745 0.6164 1.2188
ELU 0.1800 1.2041 1.8703 1.5525 0.9373 0.7095 1.1177
LA-SiLU 0.1898 1.2232 1.9876 1.4006 1.1328 0.6804 1.1056
LA-HardSiLU 0.3023 1.4008 2.2405 1.7160 1.1416 0.8028 1.3129
Table 14: Standard deviation of ResNet44s’ accuracy on CIFAR10 and CIFAR10-C with different activation functions. We average in the table.
Activation Clean Total Noise Blur Digital Weather Extra
ReLU 0.6249 2.1269 3.5833 2.4709 1.8301 1.2399 1.8742
LReLU 0.4437 1.9929 3.3751 2.3684 1.6750 1.1650 1.7264
PReLU 14.5207 11.3037 8.8671 10.8069 11.8174 12.5987 11.8192
SiLU 0.2435 1.4765 2.6747 1.7969 1.0507 0.8144 1.3453
HardSiLU 0.4192 1.7214 3.2993 1.8791 1.3047 1.0844 1.4341
MISH 0.4472 2.1393 3.7470 2.5737 1.6741 1.2354 1.8683
GELU 0.4696 1.6769 3.1032 1.8889 1.2754 0.8846 1.5892
ELU 0.2073 1.4907 2.7899 1.7389 1.1132 0.7598 1.3765
LA-SiLU 0.2719 1.3634 2.5109 1.5678 1.0625 0.7445 1.2180
LA-HardSiLU 0.2033 1.5738 2.8731 1.7762 1.3225 0.8906 1.3316
Table 15: Standard deviation of ResNet20s’ accuracy on CIFAR100 and CIFAR100-C with different activation functions. We average in the table.
Activation Clean Total Noise Blur Digital Weather Extra
ReLU 0.2937 0.8120 1.0754 0.9131 0.7956 0.5861 0.7558
LReLU 0.3233 0.8633 1.0476 0.8994 0.7943 0.7862 0.8353
PReLU 0.4435 0.9101 1.2044 1.0421 0.8121 0.7970 0.7682
SiLU 0.4147 0.7665 1.0922 0.8286 0.6642 0.5197 0.8095
HardSiLU 0.3903 0.8104 1.1921 0.8172 0.6649 0.6375 0.8356
MISH 0.3269 0.7400 1.0435 0.6064 0.7377 0.6462 0.7420
GELU 0.3866 0.8059 0.9036 0.9290 0.8281 0.6220 0.7714
ELU 0.3618 0.7367 1.0898 0.7222 0.6868 0.5689 0.7044
LA-SiLU 0.3493 0.8537 1.3183 0.8409 0.7137 0.6806 0.8313
LA-HardSiLU 0.4056 0.8751 1.2555 0.8963 0.7188 0.7466 0.8536
Table 16: Standard deviation of ResNet32s’ accuracy on CIFAR100 and CIFAR100-C with different activation functions. We average in the table.
Activation Clean Total Noise Blur Digital Weather Extra
ReLU 0.5187 1.0671 1.3930 1.1459 1.0240 0.8663 0.9876
LReLU 0.3364 0.9092 1.3616 0.8028 0.9881 0.6881 0.8184
PReLU 0.5045 0.9674 1.2581 1.0727 0.8316 0.8351 0.9123
SiLU 0.5053 1.0087 1.5275 1.0608 0.8957 0.6965 0.9926
HardSiLU 0.5325 0.9114 0.9530 1.0173 0.9229 0.9203 0.7538
MISH 0.4820 0.9148 1.2750 0.9058 0.8194 0.7269 0.9369
GELU 0.4418 0.8121 1.0506 0.9158 0.7070 0.6675 0.7790
ELU 1.4877 1.3180 1.1907 1.3131 1.3876 1.4028 1.2640
LA-SiLU 0.3776 0.8765 1.3721 0.8750 0.6403 0.6864 0.9326
LA-HardSiLU 0.4486 0.9763 1.5039 1.0204 0.8062 0.7838 0.8988
Table 17: Standard deviation of ResNet44s’ accuracy on CIFAR100 and CIFAR100-C with different activation functions. We average in the table.
Activation Clean Total Noise Blur Digital Weather Extra
ReLU 0.5652 1.1011 1.5926 1.1677 0.9634 0.9098 0.9947
LReLU 0.6872 1.2173 1.8513 1.1183 1.1814 0.9895 1.1046
PReLU 0.7084 1.3226 1.8320 1.4481 1.0939 1.2049 1.1616
SiLU 0.5204 0.9563 1.2746 0.9791 0.8475 0.8444 0.9157
HardSiLU 0.9417 1.4808 1.6740 1.5590 1.5011 1.3818 1.3362
MISH 0.5119 0.9386 1.3795 0.9206 0.7863 0.7879 0.9291
GELU 0.6182 0.9809 1.3545 0.9600 0.9044 0.8583 0.9207
ELU 0.4118 1.1195 1.6396 1.1471 0.9781 0.8948 1.0679
LA-SiLU 0.4366 1.1648 1.5511 1.2528 1.1577 0.9353 1.0238
LA-HardSiLU 0.7955 1.8629 2.9063 1.8458 1.6924 1.5680 1.5630
Table 11 illustrates the classification performance of ResNet32 and ResNet44 on CIFAR10-C and CIFAR100-C datasets. We do not report the results for ResNet44 with PReLU on CIFAR10 due to network instability during training. LayerAct functions outperformed the element-level activation functions in all but one instance, specifically ResNet44 on CIFAR10-C. However, it is noteworthy that LA-SiLU and LA-HardSiLU demonstrated much robust inference compared to their corresponding element-level activation functions, SiLU and HardSiLU, in all cases. Tables 12, 13, 14, 15, 16, and 17 demonstrate the standard deviation of the networks on the CIFAR and CIFAR-C dataset.
Table 18: Statistical significance test of LA-SiLU on CIFAR10 dataset.
LA-SiLU
ReLU LReLU PReLU SiLU HardSiLU Mish GELU ELU
RN20 $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ 0.002 $<1e^{-3}$ 0.011 0.016 $<1e^{-3}$
RN32 0.02 0.011 0.005 0.331 $<1e^{-3}$ 0.422 0.125 $<1e^{-3}$
RN44 0.014 0.001 - 0.007 $<1e^{-3}$ 0.479 0.094 $<1e^{-3}$
Table 19: Statistical significance test of LA-SiLU on CIFAR100 dataset.
LA-SiLU
ReLU LReLU PReLU SiLU HardSiLU Mish GELU ELU
RN20 $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ 0.065
RN32 $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$
RN44 0.357 0.244 0.008 0.006 $<1e^{-3}$ 0.476 0.207 $<1e^{-3}$
Table 20: Statistical significance test of LayerAct functions on CIFAR10-C dataset.
LA-SiLU
ReLU LReLU PReLU SiLU HardSiLU Mish GELU ELU
RN20 $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$
RN32 $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$
RN44 0.399 0.233 - $<1e^{-3}$ 0.001 0.002 0.003 $<1e^{-3}$
LA-HardSiLU
ReLU LReLU PReLU SiLU HardSiLU Mish GELU ELU
RN20 $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$
RN32 0.007 0.014 $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$
RN44 0.18 0.138 - $<1e^{-3}$ 0.008 0.01 0.018 $<1e^{-3}$
Table 21: Statistical significance test of LayerAct functions on CIFAR100-C dataset.
LA-SiLU
ReLU LReLU PReLU SiLU HardSiLU Mish GELU ELU
RN20 $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$
RN32 $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$
RN44 $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$
LA-HardSiLU
ReLU LReLU PReLU SiLU HardSiLU Mish GELU ELU
RN20 $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$
RN32 $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$
RN44 $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$ $<1e^{-3}$
[…]
Appendix K K. Additional Figures
Figure 8: Distribution of the activation input means of the elements in a trained network on MNIST at $1^{st}$ and $40^{th}$ epochs.
Figure 9: Distribution of activation input fluctuation due to noise with different noise distribution.
In this section, we present additional tables and figures extracted from the experiments. Figure 8 presents the distribution of the mean activation input. As observed in the mean of activation input at epoch 40 (right), LayerAct functions promote the training of parameter $W$ such that the output of the linear projection $y=W^{T}x$, which is also activation input, gets closer to zero compared to other functions. This helps the activation output to exhibit a ‘zero-like’ behaviour.
LayerAct functions exhibit a significantly lower mean and variance of activation fluctuation among the samples compared to any other element-level activation function (see Figure 3 in the main manuscript). Figure 9 demonstrates that the distribution of mean fluctuation in activation input appears similar across all functions. This observation confirms that the lower mean and variance of activation output fluctuation of LayerAct functions is not due to a smaller fluctuation in activation input, but is a result of the inherent mechanism of LayerAct.
Appendix L L. Medical Image
In this section, we present the setting of experimental result from U-net (Olaf Ronneberger 2015) and Unet++ (Zhou et al. 2018) for segmentation task on a nuclei image dataset from Data Science Bowl 2018 (Goodman et al. 2018). Detailed experimental setting is as follows: (1) Adam optimizer with $3e^{-4}$ learning rate, and $1e^{-4}$ weight decay, (2) training 100 epoches with cosine annealing scheduler, and (3) BCE-Dice Loss as the loss function. We report the average IoU (Intersection over Union; %) over 10 trials with different weight initialization in Table 4 of the main manuscript.
Appendix M M. Adversarial Robustness
We conducted experiments to investigate the adversarial robustness of LayerAct. We utilized ReLU, SiLU, and LA-SiLU as the activation of ResNet18. The experimental setting was based on (Zhou et al. 2023). The experimental results in Table 4 in the main manuscript show that the ResNet20 with LA-SiLU is more robust against adversarial attacks compared to those with other activations. These results indicate that LayerAct can enhance adversarial robustness as well.